The length of a diagonal of a rectangle is 10cm. The diagonal is inclined at an angle θ° to a side of the rectangle. Show that the perimeter,P cm, of the rectangle is given by P=20√2 sin (θ+45)°
Answer by Greenfinch(383)

P = 2x 10 sin y + 2x 10 cos y
P = 20 (sin y + cos y)
Convert to R sin (y + a) format
R = 20 sqrt( 1 + 1)= 20 sqrt 2
tan a = 1/1 =1 a = 45
So P = 2- sqrt 2 ( sin (y - 45))